OpenStudy (anonymous):

What is the general solution for r^6+16r^2=0?

OpenStudy (perl):

$$\Large{ r^6+16r^2=0? \\r^2( r^4 + 16) = 0 \\ \Rightarrow r^2 = 0 ~,~ r^4 + 16 = 0 }$$

OpenStudy (anonymous):

r^4 = -16 im stuck after that

OpenStudy (perl):

that does not have real solutions,

OpenStudy (perl):

are you looking for the imaginary solutions as well

OpenStudy (anonymous):

yeah

OpenStudy (thomas5267):

\[ r^4=-16\\ r=2\sqrt[4]{-1} \] So find the quartic root of -1.

OpenStudy (anonymous):

+/- 2i?

OpenStudy (perl):

right

OpenStudy (anonymous):

I need the answer in the form y=c1+c2... but I only know r1 and r2=0 and r4,r5,r6 are missing.

OpenStudy (perl):

this is a differential equation, can you state the original problem

OpenStudy (anonymous):

Find the general solution to y^6+16y''=0

OpenStudy (thomas5267):

I don't think the answer are -2i and 2i and (-2i)^4=(2i)^4=16.

OpenStudy (anonymous):

yeah thats what i was thinking then id still have r6 missing

OpenStudy (perl):

r^4 = -16 r^2 = ± sqrt(-16) r^2 = ± 4i r = ± √ (± 4i)

OpenStudy (anonymous):

should i use -1= e^i(pi+2npi) for that

OpenStudy (thomas5267):

\[ -1=\cos(\pi)+i\sin(\pi)\\ \sqrt[4]{-1}=\cos\left(\frac{\pi}{4}+\frac{n\pi}{2}\right)+i\sin\left(\frac{\pi}{4}+\frac{n\pi}{2}\right),\,n=0,1,2,3 \]

OpenStudy (anonymous):

gotcha thanx!

OpenStudy (perl):

r^4 = -16 r^2 = ± sqrt(-16) r^2 = ± 4i r = ± √ (± 4i) r= √( 4i), √(-4i), -√(4i), -√(-4i)

OpenStudy (perl):

r^4 = -16 r^2 = ± sqrt(-16) r^2 = ± 4i r = ± √ (± 4i) r= 2√i, 2√-i, -2√i, -2√i

OpenStudy (anonymous):

@perl would doing that thomas' way give me the same answer?

OpenStudy (perl):

sure :)

OpenStudy (anonymous):

got it thank you guys!

OpenStudy (perl):

√i = √(2)/2 + i √(2)/2

OpenStudy (perl):

Demoivre's theorem is an good approach to finding the nth root of a number

OpenStudy (thomas5267):

I have no idea what I am doing but could we try this? \[ y^{(6)}+16y''=0\\ y^{(5)}+16y'=a\\ y^{(4)}+16y=ax+b\\ \lambda^4+16=0\\ \lambda=\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}},\,-\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}},-\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}},\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}}\\ y=c_1e^{\left(\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}}\right)x}+c_2e^{\left(-\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}}\right)x}+c_3e^{\left(-\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}}\right)x}+c_4e^{\left(\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}}\right)x}+ax+b \]

OpenStudy (thomas5267):

Just in case you can't see this \[ \begin{align*} y&=c_1e^{\left(\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}}\right)x}+c_2e^{\left(-\frac{\sqrt{2}}{{2}}+i\frac{\sqrt{2}}{{2}}\right)x}\\ &\quad +c_3e^{\left(-\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}}\right)x}+c_4e^{\left(\frac{\sqrt{2}}{{2}}-i\frac{\sqrt{2}}{{2}}\right)x}\\ &\quad +ax+b \end{align*} \]

OpenStudy (anonymous):

thanx again @thomas5267 i appreciate it!